OpenStudy (anonymous):

Solve each equation or formula for the variable indicated. 1. ax-b=c, for x 2. 15x+1=y, for x

OpenStudy (anonymous):

ax= b + c x = (b+ c)/ 2 15x = y - 1 x = ( y-1) / 15

OpenStudy (anonymous):

But where did the 2 come from?

OpenStudy (anonymous):

sorry its a

OpenStudy (anonymous):

So all you have to do for the given variable is leave it on one side on its own?